Operation Summary for April 21, 2010 - April 23, 2010
Wednesday April 21
The day shift began with Operations monitoring the Tevatron (TeV) store 7753 (114.62E30), the Recycler stash (263.2E10), and the Antiproton Source (Pbar) stack (11.2E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.
During the evening shift, Operations established store 7754 at 4:01 PM with an average initial luminosity of 373.2E30. A MI access ran from the end of shot setup to 7:28 PM to recover the LCW flow interlocks in MI-30. Operations reported earlier that the High Voltage expert FESS brought in to investigate the Master Substation problem could not begin his work because his test equipment was faulty; he will, hopefully, begin his troubleshooting on Thursday.
Thursday April 22
The midnight shift began with Operations monitoring store 7754 (125.86E30) and the stash (87.1E10) and the Pbar stack (7.1E10), and with NuMI and MiniBooNE off due to load switch work.
During the midnight shift, at 6:38 AM, Operations reported that a Booster kicker (MKS09) was sparking; operators bypassed the kicker, which is only used for short-batch cycles. A water group person reported at 7:13 AM that he was taking CUB column #3 offline for regeneration. Operations reported some short downtimes due to problems with a Linac RF station (LRF1) and a MI-40 quadrupole power supply.
During the day shift, Operations terminated store 7754 at 10:29 AM. Operations prepared CDF, D0, and the TeV A-E areas for accesses. Fermi beam commissioners sent 104 beam pulses to the MTA - MuCool experiment; eventually they saw beam on the first multiwire, but not beyond. Operators began the TeV turn-on and checkout at 3:24 PM. FESS reported that the High Voltage expert they brought in had successfully located and repaired the ground faults in the DC controller cabinets at the Master Substation.
During the evening shift, Operations put the TeV into shot setup at 7:30 PM, but this was cut short due to a loss of antiprotons (210E10) in the Recycler; operators investigated and called in a Recycler person to help.
Friday April 23
The midnight shift began with Operations monitoring the stash (293.3E10), with Pbar stacking (121.3E10), and with NuMI and MiniBooNE taking beam.
During the midnight shift, Recycler personnel reported, at 12:41 AM, that they had found and fixed a damper problem that was causing the instability. Operations put the TeV into shot setup at 1:42 AM and established store 7758 at 2:57 AM with an AIL of 293.7E30.
The Plans for Friday and the Weekend
The plans for today are stack, stash, and store.
Comments and Suggestions
|last modified 04/23/2010 email Fermilab|
|Security, Privacy, Legal|